OpenStudy (anonymous):

Please help!!! Write the equation of the circle that satisfies this condition: The endpoints of a diameter are at (-2,-3) and and at (4,5)

OpenStudy (anonymous):

@ParthKohli @ganeshie8 @Hero @dan

hero (hero):

If you have the endpoints of the diameter of a circle, the center of the circle can be found using the midpoint formula, correct?

OpenStudy (anonymous):

(x+2)(x-4)+(y+3)(y-5)=0

OpenStudy (anonymous):

is that x1+x2/2 and y1+y2/2?

OpenStudy (anonymous):

i got (1,1)

OpenStudy (shamim):

yes

OpenStudy (anonymous):

(x+2)(x-4)+(y+3)(y-5)=0 is the required equation

hero (hero):

Very good. Now find the radius. The radius is the length from one of the endpoints to the center point of the circle.

OpenStudy (anonymous):

if the ends of diameter have coordinates (a,b) and (c,d ) then equation of circle is (x-a)(x-c)+(y-b)(y-d)=0

OpenStudy (anonymous):

so can I use (1,1) and (4,5)? the distance between these?

OpenStudy (anonymous):

@Hero

hero (hero):

Yes, you can use the distance formula to find the distance between those two points.

hero (hero):

Afterwards you'll have the radius \(r\) and the center \((h,k)\). From there, you can insert those values in to the equation of a circle: \((x - h)^2 + (y - k)^2 = r^2\)

OpenStudy (anonymous):

so radius is 5? or is it 5 squared?

hero (hero):

You calculated r already? Can you show please the work you did to calculate \(r\) if you don't mind?

OpenStudy (anonymous):



hero (hero):

That looks good. So what do you believe is the equation of the circle?

OpenStudy (anonymous):



hero (hero):

The the left side is equal to \(r^2\) not just \(r\).

OpenStudy (anonymous):

so its 25

hero (hero):

So the equation of the circle is...?

OpenStudy (anonymous):

\[(x-1)^{2}+(y-1)^{2}=25\]

hero (hero):

Excellent work.

OpenStudy (anonymous):

Thank you for your help!!!